OpenStudy (anonymous):

Write the domain and range for each function below Please 9.(5,0),(3,-2)(1,4),(-1,-6)

OpenStudy (anonymous):

Help please i also need help with functions

OpenStudy (vshiroky):

The domain is x and range is y

OpenStudy (anonymous):

so i write all the x's and all the ys?

OpenStudy (vshiroky):

For the domain you write the X unless there is 2 of the same number. So domain is {5,3,1,-1} and the range would be?

OpenStudy (anonymous):

0,2,-6

OpenStudy (anonymous):

0,2,4,-6

OpenStudy (vshiroky):

Not quite... all the y points

OpenStudy (anonymous):

Which ones would it be :p

OpenStudy (vshiroky):

0,-2,4,-6

OpenStudy (anonymous):

0,-2,4,-6

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Are you familiar with functions aswell?

OpenStudy (vshiroky):

Like?

OpenStudy (anonymous):

Let me screenshot one second

OpenStudy (anonymous):

Directions Calculate f ( x ) for the given domain values.

OpenStudy (anonymous):

Calculate f(x) for the given domain values

OpenStudy (vshiroky):

Did they give you an equation?

OpenStudy (anonymous):

Yes let me attach the file

OpenStudy (anonymous):

Page 13

OpenStudy (anonymous):

im suppose to do all of page 13

OpenStudy (anonymous):

Maybe you could help with a few of them? please

OpenStudy (vshiroky):

f(x)=5x; x 4, 6, 8, 10, 20 Just plug in x f(x)=5(4) f(x)=5(6) f(x)=5(8) f(x)=5(10) f(x)= 5(20)

OpenStudy (anonymous):

so i just put x in but how many answers would i have?

OpenStudy (vshiroky):

Every one of those equations give you an answer

OpenStudy (anonymous):

number 1 correct 5(4) =20 5(6)=30 5(8)=40 5(10)=50 5(20)=100?

OpenStudy (vshiroky):

Yep

OpenStudy (anonymous):

okay I get it thank you :)

OpenStudy (vshiroky):

Np

OpenStudy (anonymous):

PAGE 17: 1, 5, 9, 15 can you help with these aswell please.

OpenStudy (vshiroky):

I'm doing my own homework at the same time so I may not be able to get to it quickly

OpenStudy (anonymous):

I understand Thank you just when you get the chance please

OpenStudy (vshiroky):

2x+4y=8 You subtract 2x from both sides giving you 4y=-2x+8 then divide both sides by 4 y=-2/4x+2 or simplified y=-1/2x+2

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Page 18 Please when you get the chance I have no clue